OpenStudy (didierr):

The decimal 0.0046 is equivalent to what percent?

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 0.05 }\) would be, \(\large\color{#000000 }{ \displaystyle 0.05\times 100 }\) percent And that is \(\large\color{#000000 }{ \displaystyle 5 }\) percent

OpenStudy (didierr):

0.000046% 0.046% 0.46% 46%

OpenStudy (solomonzelman):

So in order to convert any number \(x\) into percent, multiply \(x\) times 100 (like I did below) \(x\times 100\) and then put a percent sign %, or the word "percent" to indicate that your answer is in percent.

OpenStudy (anonymous):

It would be .46 percent.

OpenStudy (anonymous):

46%

OpenStudy (anonymous):

Because you times decimal by 100 to get percent.

OpenStudy (sleepyjess):

Great explanation Solomon!

OpenStudy (solomonzelman):

tnx sleepujess ;)

OpenStudy (anonymous):

For Example: